Project

General

Profile

05-Dec-2019

Agenda:
1. General status updates/Schedule review.
John -- 4 STM cars (old timing cards) in PREP. Going to put fiber cassettes in PREP.

Rick -- working on reproducing problem that Calo team saw in Pisa, can reproduce with ots in Rick’s office. Tracing down why communication stops. For CRV, had email from Sten - internal CRV issues, want to get copper sync signals implemented. Sten is working on firmware.

Karen -- integration team would like to see clock interface.

Ryan -- Tracker has synchronized clock and sent Data Payload Packets (O100) this week.

Andrei - Debugging ExtMon hardware, Xilinx, 4 cards total - ¾ works.

Micol - Working with Monica, filling memory for DDR interface. FIFO getting and staying full.

Ron

Antonio -- Working on JS event handling in GUI, and resize handling. This week working on alarm propagation (from EPICS) - and ots state machine handling - seems to be functional.

Glenn -- Minor and major alarms exist in EPICS (specified by lolo, lo, hi, hihi.. For analog types; Different for different types.). Plan is to setup central db that would mimic mu2e building db. Had meeting with SL support group to specify EPICS db requirements - support group was asking 5 days or 7 days, e.g.. Service Level Agreement SLA with Microboone for computers, includes EPICS db server -- not SLA for servicing problems, but they do it anyway. Copy Microboone agreement?

Agreement that we want to configure which PV alarms stop the state machine. Plan to configure for each FSM transition and running workloop.

Vivian -- Need SLA? More disks on EPICS server? Glenn confirms 15GB disk space only.

Eric -- Start putting together list of servers, and data logger, .. for DCS02 has 733 GB free on dcs free on /scratch .. SLAM group does not mount all disks by default. Dcs01 nobody can login. Eric will open ticket.

Lei and Craig -- Called Sten this morning, still working. Sten coming this afternoon.

Greg -- Pointing out STM and ExtMon processing power is outside TDAQ scope. Conisder processing a pool - and show a few scenarios.

Nam -- STM Processing, Jim planning to use TDAQ server 1 or 2 cores.

Ryan -- posted docdb 29994.
2. Status of DTC to CRV ROC interface.
3. Status of EPICS on DAQ cluster.
4. Status of core ROC firmware testing.